A voltaic cell consists of a copper electrode in a solution of copper(II) ionsand a palladium electrode in a solution of palladium(II) ions. The palladiumis the cathode and its reduction potential is 0.951 V.

(a) Write the half-reaction that occurs at the anode.
(b) If E° is 0.609 V, what is the potentialfor the oxidation half-reaction?
(c) What is Keqfor this reaction?
1
Expert's answer
2019-10-09T07:35:13-0400

a)     Cu(s) -> Cu2+(aq) + 2e-

b)     E°cell = E°red + E°ox

0.609 V = 0.951 V + E°ox

ox = -0.340 V

-0.340 V

c)     log Keq = n * E°/0.0592

log Keq = 2 * 0.609 V/0.0592 = 20.6

Keq = 1020.6 = 3.98 * 1020

3.98 * 1020